# Transformer representation learning is necessary for dynamic multi-modal physiological data on small-cohort patients

### v2

@@ Abstract
Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing University]National Laboratory of Solid State Microstructure, Department of Physics, Nanjing University, Nanjing 210093, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China
@@ 3 Discussion
Logistic Regression and SVM alone fail in effectively distinguish between POD and non-POD patients for TYPE II. Experiments conducted on patient TYPE II reflect that standalone LR and SVM classifiers are incapable of differentiating POD patients from non-POD patients. The extremely low sensitivity suggests that distinctions between POD and non-POD patients are not well manifested in the raw physiological signals. This highlights the necessity of advanced representation methods to capture informative patterns in such challenging patient cohorts.

### v3

@@ Abstract
Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing University]National Laboratory of Solid State Microstructure, Department of Physics, Nanjing University, Nanjing 210093, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China
@@ 3 Discussion
Logistic Regression and SVM alone fail in effectively distinguish between POD and non-POD patients for TYPE II. Experiments conducted on patient TYPE II reflect that standalone LR and SVM classifiers are incapable of differentiating POD patients from non-POD patients, according to Table 7. The extremely low sensitivity suggests that distinctions between POD and non-POD patients are not well manifested in the raw physiological signals. This highlights the necessity of advanced representation methods to capture informative patterns in such challenging patient cohorts.